23. Various financial data for SunPath Manufacturing for 2012 and 2013 follow.

2015

2016

Output: Sales

$300,000

$330,000

Inputs: Labour

$40,000

$43,000

Raw Materials: $45,000

$51,000

Energy:

$10,000

$9,000

Capital

$250,000

$262,000

Employed:

Other:

$2,000

$6,000

What is the percentage change in SunPath’s total productivity measure between 2015 and 2016?

  1. -9.22
  1. 2.33
  1. -0.53

D.   2.88

E. 10.39

24. A company wants to generate a forecast for unit demand for year 2017 using exponential smoothing. The actual demand in year 2016 was 120. The forecast demand in year 2016 was 110. Using this data and a smoothing constant alpha of 0.2, which of the following is the resulting year 2017 forecast value?

  1. 100
  1. 110

C.   111

D.   112

E.   120

26. Which of the following is a possible source of bias error in forecasting?

a. Failing to include the right variables

B. Using the wrong forecasting method

C. Employing less sophisticated analysts than necessary

D. Using incorrect data

E. Using standard deviation rather than MAD

35. If the best operating level of a piece of equipment is at a rate of 400 units per hour and the actual output during an hour is 300 units, which of the following is the capacity cushion?

a. 25 percent

b. 100 units per hour

C. 75 percent

D. 125 percent

E. 133 percent
Answer #1

23)

For 2015,

Output = 300000

Total Input = 40000+45000+10000+250000+2000 = 347000

Productivity = Output/Input = 300000/347000 = 0.86455

For 2016,

Output = 330000

Total Input = 43000+51000+9000+262000+6000 = 371000

Productivity = 330000/371000 = 0.88949

Change in Productivity = (0.88949 - 0.86455)/0.86455 = 0.0288 = 2.88%

Correct Option is D

24.

Forecast for 2017 = Forecast for 2016 + smoothing constant(Actual of 2016 - Forecast for 2016)

Forecast for 2017 = 110 + 0.2(120-110) = 110 + 2 = 112

Correct Option is D

26

Correct option is A

27

Capacity Cushion = Actual Output/Best operational capacity = 300/400 = 0.75

Correct Option is C.